OpenStudy (anonymous):

If 45.6 grams of water decomposes at 301 Kelvin and 1.24 atmospheres, how many liters of oxygen gas can be produced? Show all of the work used to solve this problem. 2 H2O (l) yields 2 H2 (g) + O2 (g) Hey I just wanted to know if any of you guys can help me figure out this question please and not just the answer it would be great thank you.

OpenStudy (dtan5457):

You need the ideal gas law here, which is pV=NRT p=pressure v=volume n=# of moles t=temperature r=molar gas content (0.0821) You get the amount of moles there are for oxygen based on h2o. 2h2o=36 grams 45.6/36=1.27 moles there are 1.27 moles of oxygen now just plug everything in

OpenStudy (dtan5457):

remember you are solving for V

OpenStudy (dtan5457):

V = nRT / P = (1.27)(0.0821)(301) / (1.24) = 25.2 L O2

OpenStudy (anonymous):

oh so that would be the step by step process to do ths type of question? thanks this really helped a lot!